it is assumed that labour productivity of particular members in a factory can be described by a normal distribution with sigma=3.what is probability that a sample mean calculated for 9 randomly chosen workers differs from the population mean by more than 2 units?

Answer by stanbon(48516)

it is assumed that labour productivity of particular members in a factory can be described by a normal distribution with sigma=3.what is probability that a sample mean calculated for 9 randomly chosen workers differs from the population mean by more than 2 units?
-----------------------------------
The std of the sample means is 3/sqrt(3).
-
Comment: You did not post a mean value for the population
and I made the mistake of assuming it was 3. All the
work I did is based on that assumption.
Check your problem statement. Maybe that phrase
"more that 2 units" refers to standard deviation
units.
-------------------------------------------------------
If x-bar differs from the mean by more that 2 units,
x-bar is less than 1 or x-bar is greater than 5.
------
z(1) = (1-3)/sqrt(9) = -2/3
z(5) = (5-3)/sqrt(9) = 2/3
-----
P(x-bar < 1)= P(z < -2/3) = 0.2525
P(x-bar > 4)= P(z > 2/3) = 0.2525
----
Ans: 0.2525+0.2525 = 0.5050
========================================
Cheers,
Stan H.